can some one check this for me please?
i have a triangle and i have to answer in simplified radical form.
20cm is on the left 15 cm is on the bottom and x is on the right.
x^2+15cm^2=10cm^2
=x^2+225cm=100cm
x^2=125
x=125
x=5sqrt25

Answer by stanbon(75887)

i have a triangle and i have to answer in simplified radical form.
20cm is on the left 15 cm is on the bottom and x (the hypotenuse)
is on the right.
------------
Using Pythagoras you get:
x^2=15^2+20^2
x^2=225+400
x^2=625
x=25 cm
---------------
Cheers,
Stan H.